# Compositional simulation-based inference for time series

### v1

@@ 4.4 Kolmogorov flow
where $\mathbf{u}$ is the velocity field, Re is the Reynolds number, $\rho$ is the fluid density, $p$ is the pressure field, and $\mathbf{f}$ is the external forcing. Following Kochkov et al. (2021); Rozet & Louppe (2023), we consider a two-dimensional domain $[0,2\pi]^{2}$ with periodic boundary conditions and an external forcing $\mathbf{f}$ that corresponds to Kolmogorov forcing with linear damping Chandler & Kerswell (2013); Boffetta & Ecke (2012). We take the Reynolds number and density as free parameters ${\bm{\theta}}=(\text{Re},\rho)$. We utilize the jax-cfd library (Kochkov et al., 2021) to solve the Navier-Stokes equations on a $64\times 64$ grid.
In contrast to Rozet & Louppe (2023), we consider the problem of parameter inference on ${\bm{\theta}}$. We apply FNSE to this problem, using only 200k transition evaluations, which successfully recovers an amortized posterior estimator that generalizes to long-term observations (Fig. 4). Notably, in such a high-dimensional state space, a simple proposal distribution would be ineffective. But, we can use both the initial distribution and simulator to propose a variety of feasible states (details in Appendix Sec. C.4). The posterior distributions for observations with both $T=10$ and $T=100$ are well-concentrated around the true parameter values that generated the specific observations (Fig. 4a,b). Posterior predictive samples are, on average, significantly closer to the observed data compared to prior predictives (in mean absolute error calculated over 1000 predictive samples from 50 different observations). Although the difference is less pronounced for $T=100$, this is primarily due to the divergence for slight parameter modification on long simulations. To quantitatively further investigate the performance, we perform a simulation-based calibration analysis (Talts et al., 2020) (Fig. 9). Overall, given the constraints on the simulation budget, the posterior calibration is good but deteriorates for larger time series.
Fig 4: Kolmogorov flow experiment: Top row shows a single selected observation. We visualize the posterior distribution considering (a, left top) only using the first ten transitions of the selected observation, along with a quantitative comparison of the mean absolute error (MAE) between posterior predictive samples and prior predictive samples of ten different observations (left, bottom), including the one shown above. The vorticity of two selected predictive samples is visualized on the right. This analysis is then repeated for the entire observation (b).
@@ 5.1 Related Work
We build upon previous work in the i.i.d. observation setting (Geffner et al., 2023; Linhart et al., 2024; Boelts et al., 2022), extending these methods to Markovian time series. In the context of data assimilation, which aims to estimate $p({\bm{x}}^{0:T})$, Rozet & Louppe (2023) introduced a technique for learning a local score network that estimates the score of $p({\bm{x}}^{t})$ based on a limited neighborhood ${\bm{x}}^{t-k:t+k}$ (i.e., its Markov blanket). In contrast, our approach emphasizes identifying a local inverse problem that can be composed to produce a global solution. Recently, a simulation-based conditional kernel density approximation for inference in SDEs was proposed (Cai et al., 2023), which can be viewed as a special case of FNLE, substituting normalizing flows with kernel density estimation.
Moreover, any Kalman filtering method (Kalman, 1960; Wan & Van Der Merwe, 2000; Arasaratnam & Haykin, 2009; Julier & Uhlmann, 1997) is inherently related, as these methods can approximate the marginal log-likelihood in an online (non-amortized) manner, but do require the availability of likelihoods. Finally, parameter inference for (stochastic) differential equations has been extensively studied across various fields, including numerical analysis (Plessix, 2006) and probability theory (Brouste et al., 2014), both of which rely on the explicit likelihood functions.
@@ 5.2 Limitations and Future Work
In this study, our primary focus is on time-invariant Markovian stochastic processes, with an extension to encompass time-varying cases and processes featuring parameterized initial distributions. We note that Hidden Markov Models, although widely applied in practical scenarios, do not fall within the scope of our method. Various approaches are viable, particularly in combination with data assimilation techniques (Rozet & Louppe, 2023) the Markovian hidden state is recovered to which our technique can be directly applied. Importantly, our approach applies to a broad spectrum of probabilistic models: first, identify a smaller inference problem from which the global target can be recovered; second, locally solve it. However, addressing these aspects remains a direction for future research.
The score corrections caused by score aggregation described in Subsection 3.2.2 could be further improved. In their methods, Linhart et al. (2024) approximates the data distribution by a single Gaussian distribution, which can be violated in real-world scenarios. Our results demonstrate that this approximation empirically still performs well even if these assumptions are strongly violated.
@@ 5.3 Conclusions
We introduced a simulation-efficient framework for amortized inference in Markovian simulators. Although flexible embedding networks are commonly used for high-dimensional time series (Lueckmann et al., 2017; Radev et al., 2020; Schmitt et al., 2023), they often demand extensive amounts of simulations and can be fragile when faced with data perturbations (Cannon et al., 2022; Gloeckler et al., 2023). Success in these methods hinges on the embedding network’s ability to capture robust and generalizable representations. In contrast, our approach decomposes the inference task into smaller, locally solvable problems, reducing computational costs and enhancing scalability for large-scale, complex simulations.
@@ Acknowledgments
5.4 Software and data
We use JAX (Bradbury et al., 2018) as the computational backbone and hydra (Yadan, 2019) to track configuration. For reference implementation of baselines, we use sbi (Tejero-Cantero et al., 2020a). Code to reproduce the experiments can be found in https://github.com/mackelab/markovsbi/.
Author contributions
@@ A.1 Derivation of Score Factorization (3)
It concludes the proof. Here, the final equality is derived from the the assumption $p({\bm{x}}^{0}|{\bm{\theta}})=p({\bm{x}}^{0})$.
@@ A.2 Extensions
A.2.1 Time Variant Simulation
The assumption of time-invariant transition where $p({\bm{x}}^{t+1}|{\bm{x}}^{t+1},{\bm{\theta}})$ does not depend on $t$ in the main manuscript can be relaxed. Regarding FSNE, the score decomposition (3) also holds for a time-variant simulator, but the score $\nabla_{\bm{\theta}}\log\tilde{p}({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$ depends on time $t$. To address the issue, we train a time-variant score $s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1},t)$ using training data ${\mathcal{D}}=\left\{({\bm{\theta}}_{i},{\bm{x}}_{i}^{t_{i}},{\bm{x}}_{i}^{t_{i}},t_{i})\right\}$, where $t_{i}$ is drawn from a random number generator, ${\bm{x}}_{i}^{t_{i}}$ is drawn from a proposal distribution $\tilde{p}({\bm{x}})$, and ${\bm{x}}^{t_{i}+1}_{i}\sim\mathcal{T}({\bm{x}}^{t_{i}+1}|{\bm{x}}^{t_{i}}_{i},{\bm{\theta}}_{i})$. FNPE and FNRE can be handled in the same way as FNLE by adding $t$ as a variable. Experimental evaluations are presented in Appendix C.1.
@@ A.2.4 Partial Factorization
Instead of considering just a single transition, the methodology naturally extends to multiple transitions, as also explored by Geffner et al. (2023). In PFNSE with $M$, we target $\nabla_{\bm{\theta}}\log p({\bm{\theta}}|{\bm{x}}^{t},\dots,{\bm{x}}^{t+M})$, while in PFNLE or PFNRE, we target $p({\bm{x}}^{t},\dots,{\bm{x}}^{t+M}|{\bm{\theta}})$ (up to constants), using a dataset of $M$-step simulations. A key advantage of partially factorized methods is that they require fewer network evaluations for inference on a fixed-size observation, which can accelerate sampling and reduce the accumulation of local errors over time. However, they may require more simulations for effective training.
Appendix B Experiment Detail
B.1 Tasks

### v2

@@ 4.4 Kolmogorov flow
Fig 4: Kolmogorov flow experiment: Single example observation of length hundred (top row). We visualize the posterior distribution condition on the observation up to $t=10$ (a, left top), along with a quantitative comparison of the mean absolute error (MAE) between posterior predictive samples and prior predictive samples on fifty different observations (left, bottom), including the one shown above. The vorticity of two selected predictive samples is visualized on the right. This analysis is then repeated for the entire observation (b).
@@
where $\mathbf{u}$ is the velocity field, Re is the Reynolds number, $\rho$ is the fluid density, $p$ is the pressure field, and $\mathbf{f}$ is the external forcing. We added Gaussian noise with standard deviation $\sigma=5\cdot 10^{-3}$ after each transition. Following Kochkov et al. (2021); Rozet & Louppe (2023), we consider a two-dimensional domain $[0,2\pi]^{2}$ with periodic boundary conditions and an external forcing $\mathbf{f}$ that corresponds to Kolmogorov forcing with linear damping Chandler & Kerswell (2013); Boffetta & Ecke (2012). We take the Reynolds number and density as free parameters ${\bm{\theta}}=(\text{Re},\rho)$.
In contrast to Rozet & Louppe (2023), we consider the problem of parameter inference on ${\bm{\theta}}$. We apply FNSE to this problem, using only 200k transition evaluations, which successfully recovers an amortized posterior estimator that generalizes to long-term observations (Fig. 4). Notably, in such a high-dimensional state space, a simple proposal distribution would be ineffective. But, we can use both the initial distribution and simulator to propose a variety of feasible states (details in Appendix Sec. E.5). The posterior distributions for observations with both $T=10$ and $T=100$ are well-concentrated around the true parameter values that generated the specific observations (Fig. 4a,b). Posterior predictive samples are, on average, significantly closer to the observed data compared to prior predictives (in mean absolute error calculated over 1000 predictive samples from 50 different observations). Although the difference is less pronounced for $T=100$, this is primarily due to the divergence for slight parameter modification on long simulations. To further investigate the performance, we perform a simulation-based calibration analysis (Talts et al., 2020) (Fig. 12). Overall, given the constraints on the simulation budget, the posterior calibration is okay but deteriorates for larger time series, indicating more simulations might be necessary to improve performance on $T=100$.
@@ 5.1 Related Work
We build upon previous work in the i.i.d. observation setting (Geffner et al., 2023; Linhart et al., 2024; Boelts et al., 2022), extending these methods to Markovian time series. In the context of data assimilation, Rozet & Louppe (2023) introduced a method to learn a local score network to estimate $p({\bm{x}}^{0:T})$ using a neighborhood ${\bm{x}}^{t-k:t+k}$. In contrast, we focus on solving a local inverse problem that leads to a global solution. Other work aims to incorporate the task structure (approximately) into the estimating neural network architecture (Weilbach et al., 2020; 2023; Gloeckler et al., 2024a). A different line of work aims to tackle high-dimensional SBI using summary statistics. These can be handcrafted  (Alsing & Wandelt, 2018), learned (Chen et al., 2021; 2023) or based on path signatures (Dyer et al., 2021a; 2022). However, estimating statistics that generalize across sequence lengths from sparse training can be challenging. We enhance the simulation efficiency of amortized SBI by leveraging the Markovian structure, which is different from sequential training schemes (Glöckler et al., 2022; Sharrock et al., 2024; Greenberg et al., 2019; Durkan et al., 2020; Deistler et al., 2022; Gutmann et al., 2016; Warne et al., 2022), which adaptively make simulations more informative for specific observations (but are thus not amortized). Similarly, other lines of work aim to distribute simulations based on a cost function (Bharti et al., 2024) and increase simulation diversity (Niu et al., 2023). The simulation efficiency of our approach could be further enhanced by incorporating such techniques. Recently, a simulation-based conditional kernel density approximation for inference in SDEs was proposed (Cai et al., 2023), which can be viewed as a special case of FNLE, substituting normalizing flows with kernel density estimation. Efficient inference in time series models was also previously approached in the ABC setting (Warne et al., 2022; Prangle, 2016; Bernton et al., 2019; Dyer et al., 2021b).
Moreover, any Kalman filtering method (Kalman, 1960; Wan & Van Der Merwe, 2000; Arasaratnam & Haykin, 2009; Julier & Uhlmann, 1997) is inherently related to FNLE, as these methods can approximate the marginal log-likelihood in an online (non-amortized) manner, but do require the availability of likelihoods (and are used to perform inference e.g. (Brouste et al., 2014)). The field of amortized inference with tractable likelihoods is inherently related (Le et al., 2017; Choi et al., 2019; Ganguly et al., 2023; Margossian & Blei, 2024).
@@ 5.2 Limitations and Future Work
Our primary focus here was on inference for Markovian stochastic processes that share the same transition distribution across time. We also sketched extensions to cases where it varies with time, as well as to processes with parameterized initial distributions. We note that Hidden Markov Models, although widely applied in practical scenarios, do not fall within the scope of our method. In combination with data assimilation techniques (Rozet & Louppe, 2023) the Markovian hidden state can be recovered to which our technique can be directly applied. Extending our approach to more general probabilistic models remains a direction for future research.
The score corrections caused by score aggregation described in Subsection 3.2.2 could be further improved. In their methods, Linhart et al. (2024) approximates the data distribution by a single Gaussian distribution, which can be violated in real-world scenarios. Our results demonstrate that this approximation empirically still performs well even if these assumptions are violated.
@@ 5.3 Conclusions
We introduced a simulation-efficient approach for amortized inference in Markovian simulators. Although flexible embedding networks are commonly used for high-dimensional time series (Lueckmann et al., 2017; Radev et al., 2020; Schmitt et al., 2023), they often demand extensive amounts of simulations and can be fragile when faced with data perturbations (Cannon et al., 2022; Gloeckler et al., 2023). Success in these methods hinges on the embedding network’s ability to capture robust and generalizable representations. In contrast, our approach decomposes the inference task into smaller, locally solvable problems, reducing computational costs and enhancing scalability for large-scale, complex simulations.
Software and data
We use JAX (Bradbury et al., 2018) as the computational backbone and hydra (Yadan, 2019) to track configurations. For reference implementation of baselines, we use sbi (Tejero-Cantero et al., 2020a; Boelts et al., 2024). Code to reproduce the experiments can be found in https://github.com/mackelab/markovsbi/.
Author contributions
@@ A.1 Derivation of Score Factorization (3)
Which concludes the proof. Here, the final equality is derived from the assumption $p({\bm{x}}^{0}|{\bm{\theta}})=p({\bm{x}}^{0})$.
@@ A.2 Extensions
A.2.1 Time inhomogeneous Simulation
The assumption that $p({\bm{x}}^{t+1}|{\bm{x}}^{t+1},{\bm{\theta}})$ does not depend on $t$, referred to as time-homogeneous transition simulations, in the main manuscript can be relaxed. Regarding FSNE, the score decomposition (3) also holds for a time-inhomogeneous simulators where $p({\bm{x}}^{t+1}|{\bm{x}}^{t+1},{\bm{\theta}})$ varies among the time; the score $\nabla_{\bm{\theta}}\log\tilde{p}({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$, however, then depends on time $t$. To address the issue, we train a time-inhomogeneous score $s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1},t)$ using training data ${\mathcal{D}}=\left\{({\bm{\theta}}_{i},{\bm{x}}_{i}^{t_{i}},{\bm{x}}_{i}^{t_{i}},t_{i})\right\}$, where $t_{i}$ is drawn from a random number generator, ${\bm{x}}_{i}^{t_{i}}$ is drawn from a proposal distribution $\tilde{p}({\bm{x}})$, and ${\bm{x}}^{t_{i}+1}_{i}\sim\mathcal{T}({\bm{x}}^{t_{i}+1}|{\bm{x}}^{t_{i}}_{i},{\bm{\theta}}_{i})$. FNPE and FNRE can be handled in the same way as FNLE by adding $t$ as a variable. Experimental evaluations are presented in Appendix E.6.
@@ A.2.4 Partial Factorization
Instead of considering just a single transition, the methodology naturally extends to multiple transitions, as also explored by Geffner et al. (2023). In PFNSE with $M$, we target $\nabla_{\bm{\theta}}\log p({\bm{\theta}}|{\bm{x}}^{t},\dots,{\bm{x}}^{t+M})$, while in PFNLE or PFNRE, we target $p({\bm{x}}^{t},\dots,{\bm{x}}^{t+M}|{\bm{\theta}})$ (up to constants), using a dataset of $M$-step simulations. A key advantage of partially factorized methods is that they require fewer network evaluations for inference on a fixed-size observation, which can accelerate sampling and reduce the accumulation of local errors over time. Moreover, longer training simulations potentially lead to superior performance in long-term predictions as fewer local errors can accumulate. However, they may require more simulations for effective training in return for these desirable properties.
A.2.5 Missing Values in Time Series Observation
Our method is capable of handling missing data. The proposed factorization leverages the Markov properties of time series simulators at observation points, which remain intact even in the presence of missing data. To preserve the amortized property regarding missing points, we can include the time step size $\Delta t$ as an additional input variable to the model. This allows to skip over timepoints that are missing and generally allows for irregular time grids.
Appendix B Score composition rules
[…]
B.1 Method by Geffner et al. (2023)
Geffner et al. (2023) avoids the issue by constructing an alternative sequence of distributions $q_{a}~{}(0\leq a\leq A)$ that satisfies $q_{0}=p({\bm{\theta}}|{\bm{x}}^{0:T})$ and $q_{A}=\mathcal{N}({\bm{\theta}};\mu_{A},\frac{\sigma_{A}^{2}}{T}I)$. Adapting their proposal to the Markovian case, we obtain
While this approach is generally applicable, it renders the backward SDE (Eq. 2) invalid as the marginals no longer follow the corresponding forward diffusion process. This requires the use of MCMC-based sampling corrections (in the sense of predictor-correct diffusion samplers (Song et al., 2021)), such as unadjusted Langevin dynamics (Geffner et al., 2023) or similar approaches (Sjöberg et al., 2024).
B.2 Method by Linhart et al. (2024)
Linhart et al. (2024) have recently proposed an alternative approach by approximating the marginal scores derived from the forward diffusion process. This enables the utilization of the reverse SDE and consequently allows for the application of standard diffusion-based sampling techniques. They showed that the correct score could indeed be written as suggested in Eq. 3 plus an additional but intractable correction term of the form
$\nabla_{\bm{\theta}}\ell({\bm{\theta}},{\bm{x}}^{0:T}_{o})=\log\int p({\bm{\theta}}|{\bm{\theta}}_{a})^{1-T}\prod_{t=1}^{T}p({\bm{\theta}}|{\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1})d{\bm{\theta}}$
To estimate it analytically, they employed Gaussian approximations,
$p({\bm{\theta}}|{\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1}):=\mathcal{N}({\bm{\theta}};\mathbf{\mu}_{a,t,t+1}({\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1}),\mathbf{\Sigma}_{a,t,t+1}({\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1}))$
$p({\bm{\theta}}|{\bm{\theta}}_{a}):=\mathcal{N}({\bm{\theta}};\mu_{a}({\bm{\theta}}_{a}),\mathbf{\Sigma_{a}}),$
making the additional terms tractable. Eventually, we have
where $\mathbf{\Lambda}({\bm{\theta}}_{a})=\sum_{t=0}^{T-1}\mathbf{\Sigma}_{a,t,t+1}^{-1}+(1-T)\mathbf{\Sigma}_{a}^{-1}$. This approach requires specifying two “hyperparameters”: the denoising prior covariance $\mathbf{\Sigma}_{a}$ in $p({\bm{\theta}}|{\bm{\theta}}_{a})$ and denoising posterior covariance $\mathbf{\Sigma}_{a,t,t+1}$ in $p({\bm{\theta}}|{\bm{\theta}}_{a},{\bm{x}}^{t},{\bm{x}}^{t+1})$. The denoising prior covariance is typically known analytically, along with the marginal prior score. Specifically, given a Gaussian prior with covariance $\mathbf{\Sigma}_{\bm{\theta}}$ we have $\mathbf{\Sigma_{a}}=(\mathbf{\Sigma}_{\bm{\theta}}^{-1}+s(a)^{2}/\sigma(a)^{2}{\bm{I}})^{-1}$ by Gaussian marginalization rules with the perturbation kernel $p({\bm{\theta}}_{a}|{\bm{\theta}})=\mathcal{N}({\bm{\theta}}_{a},s(a){\bm{\theta}},\sigma(a)^{2}I)$ specific to the diffusion model (i.e. see Appendix Sec. D.2).
In contrast, the denoising posterior covariance must be estimated. To address this, GAUSS assumes a Gaussian clean posterior (i.e., at $a=0$) and analytically computes the denoising posterior covariances (analogous to the prior, but with an estimate of $\mathbf{\Sigma}_{{\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1}}$). This estimate can be obtained through samples, given that we have a diffusion model that indeed can sample from $p({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$. Alternatively, JAC directly estimates the denoising covariance iteratively using Tweedie’s Moment projection (Boys et al., 2024), leveraging the Jacobian of the score network:
$\mathbf{\Sigma}_{a,{\bm{x}}^{t},{\bm{x}}^{t+1}}=\frac{m(a)^{2}}{s(a)^{2}}\left(\mathbf{I}+\sigma(a)^{2}\nabla_{{\bm{\theta}}_{a}}s_{\phi}({\bm{\theta}}_{a}|{\bm{x}}^{t},{\bm{x}}^{t+1})\right)^{-1}$
If the Gaussian assumption is satisfied, both approaches are theoretically equivalent (neglecting numerical approximation errors). Yet, if these are violated, these yield different approximations to the denoising covariance matrices and hence will behave differently.
B.3 Practical implementation details
The GAUSS and JAC approximation is only valid if $\mathbf{\Lambda}$ is positive definite, as the derivation of this form depends on it. Ignoring this can lead to numerical instabilities. Unlike Linhart et al. (2024), we do not clip the diffusion process to a specific prior range to avoid numerical issues; we adapt all estimated posterior precision matrices to ensure positive definiteness of $\mathbf{\Lambda}$. This adaptation resolves the problem even for long time series (i.e., $T>100$; see Appendix E.3 for details). Importantly, since the GAUSS method is derived under Gaussian assumptions, it presupposes that the posterior covariance matrices (and their corresponding scores) are smaller than those of the prior due to observational constraints. This assumption may not hold true, particularly in multi-modal tasks. For instance, in the Periodic SDE task, four modes are symmetrically positioned around the origin. While the variances of each mode may decrease, the total variance often still increases, especially when the modes are distant from the origin. This situation can render $\Lambda({\bm{\theta}}_{a})$ non-positive definite, thus invalidating the approximation. Similarly, for JAC, the estimated posterior covariance might not even be positive and definite as the Jacobian of the scoring network might badly represent the Jacobian of the true score.
We utilize the GAUSS composition rule as our default method: We estimate the posterior precision matrices $\mathbf{\Sigma}_{a,t,t+1}^{-1}$ as recommended by Linhart et al. (2024). For each posterior $p({\bm{\theta}}|{\bm{x}}^{t},{\bm{x}}^{t+1})$, we produce $500\cdot\dim_{\bm{\theta}}$ samples in parallel from the diffusion model. These samples are then used to estimate the associated covariance matrix. Although this approach introduces a slight constant overhead prior to each sampling run, it does not require the iterative estimation of Jacobians, as is necessary with the JAC method. On the other hand, within our settings, the dimension of the parameter is often much smaller than the data dimension, rendering computation of the Jacobian not too expensive.
Appendix C Practical guide to proposal construction
Fig 5: Proposal: A set of trajectories within the phase space of the Lotka Volterra task for different parameterizations sampled from the prior (constructed using a total of 5k simulation steps) (a). The proposal was constructed by randomly sampling noisy points from the state space trajectories (b). Posterior approximation in sliced Wasserstein distance constructed by improved proposal compared to our “naively” chosen baseline (c).
In this section, we provided a practical guideline for constructing the proposal $\tilde{p}({\bm{x}}^{t})$. In general, the proposal will be task-specific. For an amortized approximation, a valid proposal $\tilde{p}({\bm{x}}^{t})$ should satisfy the conditions: (i) It must have support at any point ${\bm{x}}^{t}$ the simulator can attain (at evaluation time). (ii) The proposed point ${\bm{x}}^{t}$ must be independent of the parameter ${\bm{\theta}}$ involved in the transition to ${\bm{x}}^{t+1}$ to ensure Eq. 3 remains valid (Appendix Sec. A.1). But solely relying on these properties does not necessarily make up a “good” proposal. Essentially, the proposal specifies which regions of the data domain are represented in the training dataset. On a finite simulation budget, we thus would expect a better posterior approximation for states that are more likely to be generated by the proposal than those that are less likely. This provides an opportunity and challenge to improve the proposal distribution design for a given simulator.
Task Periodic SDE Linear SDE Mixture RW Double Well
T=1 T=11 T=101 T=1 T=11 T=101 T=1 T=11 T=101 T=1 T=11 T=101
FNSE 0.53 0.60 0.76 0.53 0.70 0.90 0.52 0.69 0.69 0.50 0.59 0.81
FNSE (pred. prop) 0.53 0.60 0.71 0.53 0.70 0.90 0.52 0.69 0.75 0.50 0.57 0.80
Table 1: Predictive proposal: Posterior approximation for different proposals given as C2ST metric for each benchmark task. Here, we compare our default proposal (trained with 100k step simulations), chosen by domain knowledge, and the prior predictive construction (using 5k simulation steps) training on the remaining 95k step simulations.
Lotka Volterra SIR
T=1 T=11 T=101 T=1 T=11 T=101
FNSE 0.57 0.78 0.94 0.67 0.78 0.74
FNSE (pred. prop) 0.52 0.65 0.88 0.67 0.79 0.74
Table 2: Predictive proposal: Posterior approximation for different proposals given as C2ST metric for Lotka Voletrra and SIR. Here, we compare our default proposal (trained with 100k step simulations), chosen by domain knowledge, and the prior predictive construction (using 5k simulation) trained on 95k step simulations.
This interpretation makes it fairly intuitive to know what might be a good proposal and domain knowledge can be used to design a proposal from scratch. If applicants know that the dynamics are bounded (or at least likely contained) within a certain region, then one can design a proposal that roughly covers it. If the applicant knows that the simulator has stationary distributions, then these would serve as good proposals. Yet, in general, it might be hard or suboptimal to construct a proposal in this naive way. The default proposals used in our analysis are indeed simple distributions (e.g., Gaussians) chosen based on such knowledge. Yet we here also provide some more general approaches to select reasonable proposals in Sec. C.1 and Sec. C.2.
C.1 Prior predictive construction
Condition (ii) does not generally exclude the use of the simulator to construct a good proposal and we hence can use some simulation outputs to construct a good proposal. As an example, let’s consider the Lotka-Volterra task. By sampling from the prior predictive (i.e., sampling parameters from the prior and simulating data), we can visualize the trajectories ${\bm{x}}^{0:T}$ of simulation outputs within its phase space (Fig. 5a). These states are likely to encompass most states, and we can use this to construct an effective proposal for our factorized methods, e.g., by simply sampling noisy points on phase space trajectories uniformly (Fig. 5b). For example, one can:
  • (A)
    Sample a few parameters $\{{\bm{\theta}}_{i}\}_{i=1}^{N}$ from the prior $p({\bm{\theta}})$.
  • (B)
    Perform time series simulations ${\bm{x}}^{0:T}_{i}\sim p({\bm{x}}^{0:T}|{\bm{\theta}}_{i})$ for each parameters until a certain time ${\bm{x}}^{T}_{i}$.
  • (C)
    Estimate the proposal $\tilde{p}({\bm{x}}^{t})$, e.g., by randomly selecting a state ${\bm{x}}_{i}^{t}$ from $N\cdot(T+1)$ states with some noise (which corresponds to a kernel density estimate).
What condition (ii) prevents is that we cannot directly leverage transition from prior predictive simulations $({\bm{\theta}}_{i},{\bm{x}}_{i}^{0:T})$ as each transition will depend on the same parameter, and thus, each triplet $({\bm{\theta}}_{i},{\bm{x}}_{i}^{t},{\bm{x}}_{i}^{t+1})$ will not satisfy that ${\bm{x}}_{i}^{t}$ is independent of ${\bm{\theta}}_{i}$. By constructing the proposal as above, apriori, we avoid this problem. Many time series simulators are often designed to ensure that their outputs change minimally for large $T$ (Hayashi et al., 2022; Khatiwala, 2024; Ormeño & General, 2024) (a.k.a. reach stable states or stationarity), a number of simulation trajectories for not too many transitions can, thus, be sufficient to explore the phase space.
[…]
C.2 Target-based construction
We can also create the proposal $\tilde{p}({\bm{x}}^{t})$ using any sort of (real/synthetic) set of observations ${\bm{x}}^{0:T}_{o}$ accessible prior to training. For amortized inference we can proceed analogously to Sec. C.1.
If we are only interested in inference for a single observation as in a sequential SBI setting (Papamakarios & Murray, 2016; Greenberg et al., 2019; Sharrock et al., 2024; Deistler et al., 2022), we can construct the proposal using this single observation by simply proposing observed states ${\bm{x}}_{o}^{t}$ at random. The full amortization property is lost in this way (as the support is constrained), but the simulation budget can be targeted on this more specific inference problem, which can lead to improved performance.
Appendix D Experiment Detail
D.1 Tasks